a 5 ft tall girl stands on level ground the sun is 25 degrees above the horizon. how long is her shadow?
Answer #1

Draw a picture of the situation. The height of the girl is the upright leg of a right triangle. The ray of light from the sun touching the top of her head is the hypotenuse and goes down and away from her at a 25° angle to the horizontal. Basic trig will show that

H/Ls = tan25° or

Ls = H/tan25° = 5/tan25° = 10.72 ft